What is the probability that exactly 4 customers will arrive in 1 hour, when the mean arrival rate is 3 customers per hour, and interarrival times are exponentially distributed?

a 0.3528
b 0.1847
c 0.1680
d 0.8153
e 0.6472
Answer #1

Mean arrival rate (\lambda) = 3 customers per hour

Prob(X=x) = \frac{\lambda ^{x}* e^{-\lambda }}{x!}

To find the probability that exactly 4 customers will arrive in 1 hour, we put x = 4 and \lambda = 3 in the above equation.

Prob(X=4) = \frac{\lambda ^{4}* e^{-\lambda }}{4!}

= \frac{3^{4}* e^{-3}}{4*3*2*1}

= 0.1680

Hence the correct answer is option C.